### v1

# Bubble dynamics in a cavitating venturi

Abstract
Cryogenic fluids have extensive applications as fuel for launch vehicles in space applications and research. The physics of cryogenic flows are highly complex due to the sensitive nature of phase transformation from liquid to bubbly liquid and vapor, eventually resulting in cavitating flows at the ambient temperature owing to the very low boiling point of cryogenic fluids, which asserts us to classify such flows under multi-phase flow physics regime. This work elucidates the modeling of bubbly flow for cryogenic fluids such as liquid nitrogen in a converging-diverging venturi-like flow device known as cavitating venturi, a passive flow control metering device. The numerical works in literature are usually limited to modeling iso-thermal bubbly flows such as water devoid of involving energy equations because there is no occurrence of interface heat transfer as latent heat of vaporization of water is higher, unlike cryogenic fluids which are sensitive to phase change at ambient conditions. So, to realize an appropriate model for modeling cryogenic bubbly flows such as liquid nitrogen flow, the effect of heat transfer at the interface and convective heat transfer from the surrounding liquid to the traversing bubble needs to be included. Numerical modeling using an in-house code involving a finite-difference method The numerical results showed the importance of including the heat transport equation due to convection and at the interface of bubble-fluid as a significant source term for the bubble dynamics. The work is supported by computational simulation using a commercial CFD package for 2-dimensional simulations to predict a characterizing parameter, namely cavitation length. A limited flow visualization experiment using a high-speed camera is performed to study the cavitating zone length.
1 Introduction
A cavitating venturi is a simple passive flow and metering device that ensures a constant mass flow rate of the fluid that passes through it, even under varying downstream pressure conditions. The cavitating venturi has a converging section followed by a throat and a diverging section. In a venturi, the throat has a minimum area. Because of this reason, whenever a flow happens in a cavitating Venturi, the fluid suffers from minimum pressure near the throat. It so happens that the fluid pressure attains the vapor pressure, which results in cavities or vapor bubbles forming, and the flow begins to cavitate. Mixed behavior, such as bubble formation and bubble collapse from the throat to the diverging section, is typical of cavitating flows in the venturi. Such a two-phase flow is also identified with nucleation, bubble growth and collapse, fission, fusion of bubbles, re-entrant jet formation, and turbulent mixing of vapor bubbles.
Cavitating flow is a typical two-phase flow phenomenon that is common in mechanical systems like Turbo pumps (near inducer), Marine propellers, etc. When pumps operate at high rotational speeds, it results in minimum pressure near the inducer, which in turn causes vapor formation resulting in attached cavitation and traveling bubble cavitation, which are the usual cavitation patterns commonly observed in turbomachines. cavitation causes considerable damages like erosion on the surface of the impeller vanes of the pumps and eventually reduces its service life.
Besides serious disadvantages due to cavitation, cavitation positively impacts a passive flow device like cavitating Venturi, which is discussed in the following section. The following figure 2 shows cavitation in a typical cavitating Venturi.
Figure 1: Rocket Propulsion systems with (a). Active Flow device, and (b). Passive flow device
Cavitating Venturi is a Passive flow control and metering device that does not require any complicated flow control instrumentation like a Controlled valve and the associated control systems. Instead, a passive flow device is a simple mechanical device that works fundamentally on the critical or chocked flow principle. The following figure 1 depicts the Rocket propulsion system with active and passive fuel flow control systems.
Figure 2: Developed cavitation in a Venturi device [1]
Both active and passive flow control systems are commonly used in Rocket propulsion flow systems to supply liquid propellants from the fuel tank to the combustion chamber via fuel injectors. However, active flow control systems are complicated by signal actuating mechanisms for control valves, owing to which they lose their reliability. Instead, a Passive flow system equipped with a Cavitating Venturi in place of a flow control valve seems advantageous in ensuring the reliability of the flow system, owing to its simple operating mechanism, namely choked flow.
Cavitation plays a positive role in Cavitating Venturi used in Passive Flow Control systems, as Cavitation helps maintain a constant flow rate. This Constant flow rate is achieved because of critical flow or chocked flow phenomenon.
  • The critical flow condition is common in both Compressible &\&& Incompressible flows. In a Compressible flow, it is understood that the flow becomes simply chocked by accelerating the fluid to the local sonic velocity at the throat section of the nozzle (as downstream pressure waves cannot travel upstream).
  • For an incompressible fluid, the sonic velocity is significantly greater than that of a compressible fluid. So, in a Nozzle or Venturi, the accelerating incompressible fluid starts evaporating before reaching its sonic velocity, and vapor bubbles appear near the throat section and further downstream.
  • The so-formed vapor bubble cloud volume is large, obstructing the fluid flow at the throat. As pressure decreases, cavitation intensity increases, which causes the area of the bubble cloud to increase further. As a result, there is an increased resistance to the flow near the throat location, which ensures the flow rate remains unchanged, exhibiting a Cavitating flow condition.
  • The formed Vapour bubbles collapse back to ensure a single-phase liquid flow in the diverging section of the Cavitating Venturi because of the gradual increase in fluid Pressure along the diverging section.
Because of this attractive property, Cavitating Venturi finds application in Liquid and Cryogenic Stages of Rocket Engines, as fluctuating pressures in the Combustion chamber, which is equal to Venturi downstream pressure, will not affect the quantity of flow through Cavitating Venturi for a wide range of operating conditions. Much research has been done to develop numerical models to predict the characteristics of choked flow[2-4,9-11]. Also, many experimental works[5,7-8] have been aimed at understanding the phenomenon and behaviour of cavitation. It also appears from open literature that very few works exist in cryogenic Cavitating Venturi flows.
The behaviour of Cryogenic cavitating flows in a Venturi is very complex because of the very low Latent heat of Vaporization of the cryogenic fluid and hence requires a detailed study of flow behavior experimentally and validating the same using an appropriate mathematical model which can capture the actual physics of Cavitating flow behavior in a Venturi.
The following are the brief of the objectives put forth
  1. 1.
    Systematically study the applicability of 1D models and the predictive capabilities of a typical one-dimensional model that could be used for Cryogenic cavitating venturi. Though 1D models are not very rigorous in ultimately predicting the behavior of cavitating venturi, they could be very useful in predicting some key parameters. There have been many studies in 1D models and are gradually being improved for cavitating venturi with water as the working fluid. However, there seems to be limited published work on extending the numerical model and trying to characterize the behaviour of cryogenic cavitating venturi. Owing to the limitations in the 1D model, it naturally leads to the following second objective.
  2. 2.
    To explore the predictive capabilities of a 2D model of cryogenic cavitating venturi using commercial CFD solver ANSYS fluent, incorporating appropriate cavitation model. These are expected to give realistic estimates of the nature of cavitation and the cavitation length for a given set of operating conditions. Such models, however, require tuning with the experimental results. Hence, the third objective of experimental work was decided.
  3. 3.
    To design and develop an experimental facility with a planar venturi test section incorporating appropriate visualization features to understand flow behaviour when handling cryogenic liquid Nitrogen and obtain the cavitating length.
The work of Brennen [9] has been taken as a basis and appropriately extended with the inclusion of thermal term for cryogenic fluids. The numerical methodology for cavitation prediction has been extended further with 2D simulations using Ansys Fluent.
2 Numerical 1D model
In Brennen [9] model, the Rayleigh Plesset Equation for the isothermal case was assumed for solving the bubble dynamics. However, this bubble dynamics equation cannot be extended to the case of cryogenic liquids, as heat transfer effects are predominant in cryogenic liquids. It is advantageous to have an additional term in the same Brennen model [9] that accounts for heat transfer and extends for cryogenic cavitating venturi flows. The following section details models used for Numerical 1D studies for cryogenic cavitating flows.
2.1 Derivation of modified Rayleigh-Plesset equation
The following derivation systematically introduces the thermal term into the Bubble dynamics equation. Starting from the classical form of the Rayleigh Plesset Equation, the derivation for the modified Rayleigh Plesset equation is as follows
$\left[R\ddot{R}+\dfrac{3}{2}\dot{R}^{2}\right]=\dfrac{\left[P_{\nu}-P_{\infty}(t)\right]}{\rho_{L}}+\dfrac{P_{g0}}{\rho_{L}}\left[\dfrac{R_{0}}{R}\right]^{3\gamma}-\dfrac{2S}{\rho_{L}R}-\dfrac{4\mu}{\rho_{L}}\dfrac{\dot{R}}{R}$ (1)
The above equation is the simple Rayleigh-Plesset Equation for the isothermal case. The equation does not have an appropriate term for handling the heat interactions. In the above equation the vapour pressure inside the bubble which can be represented as $P_{v}(T_{B})$. In ordinary fluid such as water, there is no temperature difference between the bubble temperature $T_{B}$ and the surrounding liquid temperature $T_{\infty}$ because of the isothermal conditions. But this assumption is not valid for cryogenic liquids as the slightest temperature difference between $T_{B}$ and $T_{\infty}$ causes a considerable change in the vapour pressure inside the bubble. If there is a difference in temperature between the bubble and the surrounding liquid, then the temperature gradient $T_{\infty}-T_{B}$ becomes a potential source for heat transfer influencing the bubble growth.
By small algebraic manipulation of adding and subtracting $P_{\nu}(T_{\infty})$ in $\left[P_{\nu}(T_{B})-P_{\infty}(t)\right]$ it can be written as
$\left[P_{\nu}(T_{B})-P_{\infty}(t)\right]=\left[P_{\nu}(T_{\infty})-P_{\infty}(t)\right]+\left[P_{\nu}(T_{B})-P_{\nu}(T_{\infty})\right]$ (2)
Substituting equation 2 in 1 the equation becomes of this form
$\left[R\ddot{R}+\dfrac{3}{2}\dot{R}^{2}\right]+\dfrac{2S}{\rho_{L}R}+\dfrac{4\mu}{\rho_{L}}\dfrac{\dot{R}}{R}=\dfrac{P_{g0}}{\rho_{L}}\left[\dfrac{R_{0}}{R}\right]^{3\gamma}+\dfrac{\left[P_{\nu}(T_{\infty})-P_{\infty}(t)\right]}{\rho_{L}}-\dfrac{\left[P_{\nu}(T_{B})-P_{\nu}(T_{\infty})\right]}{\rho_{L}}$ (3)
The third term on the right-hand side of the equation is referred to as a thermal term, and the second term is the tension or the driving term for bubble growth. To find an expression for the thermal term, we need to look into the Heat Balance and B-factor theory [1] & [5]
The thermal effect in cryogenic liquids can be interpreted as the heat interaction between the vapour bubbles and the surrounding liquid medium. The production of a vapour volume $V_{\nu}$ requires the quantity of heat $\rho_{\nu}V_{\nu}L$ to be supplied (L is the latent heat of vaporization). This energy is given by the liquid near the vapour cavity. So, the liquid mass surrounding the vapour bubble gets cooled by supplying heat to the bubble. Steffano [1] & [5] put forth a theory, popularly known as B-factor theory, which gives a non-dimensional number for temperature drop called B-factor [1]
The volume of liquid $V_{L}$ after supplying the heat of vaporization reduces itself with its temperature by a drop $\Delta T$, and the heat balance between the liquid and vapour is given by
$\rho_{\nu}V_{\nu}L=\rho_{l}V_{l}C_{pl}\Delta T$ (4)
From the above equation, the B-factor is given by the ratio of the volume of vapor to the volume of liquid. Alternately, it is also given by the ratio of the actual temperature drop to the reference temperature drop, which is shown below
$B=\dfrac{\Delta T}{\dfrac{\rho_{\nu}L}{\rho_{l}C_{pl}}}=\dfrac{V_{\nu}}{V_{l}}$ (5)
Where,
$\Delta T^{\ast}=\dfrac{\rho_{\nu}L}{\rho_{l}C_{pl}}$ is the reference temperature drop and $\Delta T$ is the actual temperature drop.
This slight temperature drop $\Delta T$ in the liquid happens to be the reason for the difference in temperature between a vapor bubble and the liquid in the immediate vicinity, which affects the vapor pressure change inside the bubble, as the bubble temperature has now been slightly changed. If the drop in temperature is very small, then by Taylor series expansion, we can write
$\Delta P_{\nu}=\dfrac{dP_{\nu}}{dT}\Delta T$ (6)
The vapour pressure change with respect to the temperature $\dfrac{dP_{\nu}}{dT}$ is given by Clausius Clayperon relation[8]
$L=T\left[\dfrac{1}{\rho_{\nu}}-\dfrac{1}{\rho_{l}}\right]\dfrac{dP_{\nu}}{dT}$ (7)
Usually, the liquid density $\rho_{l}$ in the Clausius-Clapeyron relation is neglected, and the relation reduces to
$L=\dfrac{T_{\infty}}{\rho_{\nu}}\dfrac{dP_{\nu}}{dT}$ (8)
To estimate the temperature depression $\Delta T$, heat transfer between the liquid and bubble has to be considered. In literature, both conductive and convective modes of heat transfer were assumed based on the flow conditions, i.e., if there is no slip between liquid and bubble, then the conductive approach is sufficient. But in our case, assuming a convective mode of heat transfer is correct.
Assuming $Q$ as the heat flux which is exchanged at the interface of bubble and liquid, then we can write
$Q=hA\delta T$ (9)
As this heat supplied by the liquid is the actual heat source for vapourization and growth of the bubble, a balance of heat can be done as shown below
$h\delta T(4\pi R^{2})=\dfrac{d}{dt}\left(\dfrac{4}{3}\pi R^{3}\right)\rho_{\nu}L$ (10)
Rearranging the equation in terms of $\delta T$ gives,
$\delta T=\dfrac{\rho_{\nu}L}{h}\dot{R}$ (11)
Positive $\delta T$ means temperature decrement in liquid and temperature increment in vapour bubble. This, in turn, increases the vapour pressure inside the bubble, which in turn makes the bubble expand or grow. Similarly, it can be interpreted that negative $\delta T$ leads to shrinkage of the vapor bubble due to condensation.
Now, coming back to the thermal term as stated in equation 3, we can write
$\dfrac{\left[P_{\nu}(T_{B})-P_{\nu}(T_{\infty})\right]}{\rho_{L}}=\dfrac{\delta P_{\nu}}{\rho_{L}}$
Substituting $\delta P_{\nu}$ from equation[25] gives
$\dfrac{\left[P_{\nu}(T_{B})-P_{\nu}(T_{\infty})\right]}{\rho_{L}}=\dfrac{dP_{\nu}}{dT}\dfrac{\delta T}{\rho_{L}}$
again $\delta T$ can be substituted from equation[29] which results in an expression of thermal term as follows
$\dfrac{\left[P_{\nu}(T_{B})-P_{\nu}(T_{\infty})\right]}{\rho_{L}}=\dfrac{dP_{\nu}}{dT}\dfrac{\rho_{\nu}L}{\rho_{L}}\dfrac{\dot{R}}{h}$ (12)
After substituting equation[30] into [22] we finally get the modified Rayleigh-Plesset equation with thermal source term as follows
$\left[R\ddot{R}+\dfrac{3}{2}\dot{R}^{2}\right]+\dfrac{2S}{\rho_{L}R}+\dfrac{4\mu}{\rho_{L}}\dfrac{\dot{R}}{R}=\dfrac{P_{g0}}{\rho_{L}}\left[\dfrac{R_{0}}{R}\right]^{3\gamma}+\dfrac{\left[P_{\nu}(T_{\infty})-P_{\infty}(t)\right]}{\rho}_{L}-\dfrac{dP_{\nu}}{dT}\dfrac{\rho_{\nu}L}{\rho_{L}}\dfrac{\dot{R}}{h}$ (13)
Equation 13 is the governing differential equation for bubble dynamics applicable for non-isothermal liquids such as cryogenic liquids.
2.2 Modelling 1D cavitating nozzle / Venturi flow
Similar to Wang and Brennen Quasi-1D model [9] for Bubbly Cavitating flows through a Nozzle, a Quasi-1D model with thermal effects is attempted in this work for applicability with cryogenic liquids. The Continuity and momentum equations for bubbly two-phase flows along with closures, i.e., the volume fraction equation and the modified Rayleigh Plesset Equation along with the Bubble momentum equation for estimation of bubble velocity, were considered in this model for solving bubble radius
For modelling the heat transfer between bubble and liquid, a convective heat transfer model was taken from the existing Ranz and Marshall model [14]. The turbulence model was not incorporated in 1D numerical studies as handling such equations was tedious. Following assumptions were made in this model,
  1. 1.
    At any cross-section of the Venturi, the bubbles had a uniform size, and their number was assumed constant.
  2. 2.
    The bubble and the liquid temperatures were equal at the interface, and there was no friction or heat transfer between the venturi walls and the flowing fluid.
  3. 3.
    Liquid and vapour densities were assumed to be constant
Supposing that flow develops only in the x direction, the steady bubbly flow continuity and momentum equations that Wang and Brennen used are given as follows
$\dfrac{\partial(1-\alpha)A}{\partial t}+\dfrac{\partial(1-\alpha)Au}{\partial x}=0$
$\dfrac{\partial u}{\partial t}+u\dfrac{\partial u}{\partial x}=-\dfrac{1}{2(1-\alpha)}\dfrac{C_{p}}{\partial x}$
Where, $C_{p}(x,t)=\dfrac{(P(x,t)-P_{0})}{(0.5\rho_{l}u_{0}^{2})}$ is the fluid pressure coefficient $P(x,t)$ is the fluid pressure , $P_{0}$ is the upstream fluid pressure, $u_{0}$ is the upstream fluid velocity and $u$ is the velocity along the length of the profile. $A(x)$ corresponds to the area profile, which is the function of space coordinate, i.e., $x$ (being a 1D case). The mathematical expression for the area profile A(x) is
${A(x)=}\left\{1-\dfrac{1}{2}C_{p,min}\left[1-cos\left(\dfrac{2\pi x}{L}\right)\right]\right\}^{-\dfrac{1}{2}}$ (14)
$\alpha(x,t)=\dfrac{\dfrac{4}{3}\pi\eta R(x,t)^{3}}{\left(1+\dfrac{4}{3}\pi\eta R(x,t)^{3}\right)}$ is the non-dimensional void fraction, $R(x,t)$ is the non-dimensional bubble radius and $\eta$ is the non-dimensional population number.
For modelling the non-linear bubble dynamics, a modified Rayleigh–Plesset equation 13 was used
$\left[R\ddot{R}+\dfrac{3}{2}\dot{R}^{2}\right]+\dfrac{2S}{\rho_{L}R}+\dfrac{4\mu}{\rho_{L}}\dfrac{\dot{R}}{R}=\dfrac{P_{g0}}{\rho_{L}}\left[\dfrac{R_{0}}{R}\right]^{3\gamma}+\dfrac{\left[P_{\nu}(T_{\infty})-P_{\infty}(t)\right]}{\rho_{L}}-\dfrac{dP_{\nu}}{dT}\dfrac{\rho_{\nu}L}{\rho_{L}}\dfrac{\dot{R}}{h}$ (15)
The above equation can be represented in total derivative form by replacing $\dot{R}=\dfrac{DR}{Dt}$, which is given by
$\left[R\dfrac{D^{2}R}{Dt^{2}}+\dfrac{3}{2}\left(\dfrac{DR}{Dt}\right)^{2}\right]+\dfrac{2S}{\rho_{l}R}+\dfrac{4\nu_{l}DR}{RDt}=\dfrac{P_{g0}}{\rho_{l}}\left[\dfrac{R_{0}}{R}\right]^{3\gamma}\\ +\dfrac{\left[P_{\nu}(T_{\infty})-P(x,t)\right]}{\rho_{L}}-\dfrac{dp_{\nu}}{dT}\dfrac{\rho_{\nu}L_{e\nu}}{\rho_{L}h_{b}}\dfrac{DR}{Dt}$ (16)
In the above equation, the last term on the right hand side is the thermal term, and $h_{b}$ is the convective heat transfer coefficient which was modelled using the already existing Ranz and Marshall [14] model for the study of evaporation of droplets in spray drying which is given by
$h_{b}=\dfrac{Nu_{b}K_{l}}{2R}$
Where $Nu_{b}=2+0.6Reb^{\frac{1}{2}}Pr^{\frac{1}{3}}$ is the bubble Nusselt number, which in turn depends on bubble Reynolds number ($Re_{b}$), and Prandtl number (Pr) which is given by [8] as
$Re_{b}=2R|\nu-u|/\lambda_{l}$
$Pr=\dfrac{C_{pl}\mu_{l}}{K_{l}}$
Where $\lambda_{l}=K_{l}/\rho C_{p}$ is the thermal diffusivity and $K_{l}$ is the liquid thermal conductivity.
For solving bubble velocity, the bubble momentum equation formulated by Albagli et al. [13] was used, which is shown as follows
$\rho_{\nu}\dfrac{D\nu}{Dt}+\dfrac{1}{2}\rho_{l}\left[\dfrac{D\nu}{Dt}-\dfrac{Du}{Dt}\right]=-\dfrac{\partial P(x,t)}{\partial x}-\dfrac{3}{8}\rho_{l}C_{D}\times\dfrac{(\nu-u)|\nu-u|}{R}$
2.3 Wang and Brennen benchmark Results for Water with thermal term
A numerical 1D code using Matlab was written for solving the Brennen model for water as the fluid with the effects of the thermal term. The results were plotted for bubble radius, void fraction, coefficient of pressure, velocity of the liquid phase, and the growth contribution terms (thermal and pressure terms). The non-dimensional bubble radius, coefficient of pressure, void fraction, and non-dimensional void fraction plots are shown below in the figure 3,7, 8, 9corresponds to the Rayleigh Plesset bubble dynamics equation with the thermal effects incorporated.
Figure 3: Non-dimensional Bubble radius plot for water at $\alpha_{0}=2.5\times 10^{-6}$
The superimposed plot of Wang and Brennen [9] bubble radius, coefficient of pressure, and void fraction plot with the current numerical model(with thermal effects) plot is shown in the figure 4,5,6
Figure 4: Non-Dimensional Bubble radius plot from Wang &\&& Brennen [9] model and current model
Figure 5: Coefficient of Pressure plot from Wang &\&& Brennen [9] model and current model
Figure 6: Void fraction plot from Wang &\&& Brennen [9] model and current model
It can be inferred from the superimposed figures 4,5, 6 that there is no much variation in the bubble radius, void fraction, and coefficient fraction plot of current model and Brennen model [9]. However, deviations may be attributed to the additional thermal effect term. This indicated that including the thermal effect in the Rayleigh Plesset equation has no considerable role to play in the bubble dynamics of Water. The reason for this is the minimal contribution of Thermal term in water.
Figure 7: Coefficient of Pressure plot for water at $\alpha_{0}=2.5\times 10^{-6}$
Figure 8: Void fraction plot for water at $\alpha_{0}=2.5\times 10^{-6}$
Figure 9: Non-dimensional bubble radius plot for water at $\alpha_{0}=2.5\times 10^{-6}$
The figure 10 gives the plot of source terms in the Rayleigh-Plesset equation. The thermal source term is compared with a pressure source term. The values of the same are plotted for an initial void fraction of $\alpha_{0}=2.5\times 10^{-6}$, where the term $\alpha_{0}$ is the initial void fraction which has been arbitrarily assumed in Wang and Brennen model [9]. This initial void fraction significantly impacts the numerical prediction of bubble radius as this signifies the number of nucleation sites or microscopic voids in the liquid phase.
For getting a clear understanding of the domination of different source terms in the bubble dynamics, the right-hand side of the modified Rayleigh-Plesset equation containing such source terms was handled separately, and the magnitude of the same was plotted. i.e., all other terms except the thermal term were taken as the Pressure source term. The following figure 10,11 shows the contribution of the above-mentioned thermal term and pressure term.
Figure 10: Source term plots for water at $\alpha_{0}=2.5\times 10^{-6}$
Figure 11: Magnified view of pressure source term plots for water at $\alpha_{0}=2.5\times 10^{-6}$
It can be inferred from the plots that the contribution of the thermal term within the flow domain of the venturi is small and is of the order $10^{-1}$ against the contribution from the pressure term, which is of the order $10^{4}$. This points out to the fact that the explicit modeling of the thermal term in the modified Rayleigh Plesset Equation may not provide pronounced effects for fluids with the high latent heat of vapourization. This fact explains the close prediction of the Wang &\&& Brennen model results [9], where such a thermal term is not modeled.
Having benchmarked the model, the same model is now extended to predict cavitating flows of liquid nitrogen.
However, the present one-dimensional model is far from providing estimates for performance prediction of cavitating venturi, and more systematic work in developing proper mathematical models to predict the performance would be required. Hence, a systematic study of a simple planar 2D numerical study was attempted using ANSYS fluent, detailed in the following section.
2.4 Numerical 1D Results: for the present case
The numerical 1D simulation was done for liquid nitrogen, including the thermal effect, using the modified Rayleigh Plesset equation using MATLAB.
Bubble Radius, Coefficient of Pressure, and Void fraction were plotted for different cases. Since the numerical model is an initial value problem, the initial values were assumed for all the variables in line with Wang and Brennen [9] model. It was inferred that the initial void fraction had a considerable effect on the nature of the solution.
2.5 1D Numerical Test cases
2.5.1 Case-1
For an Initial void fraction $\alpha_{0}=2.5\times 10^{-6}$ and $C_{p}$=0
Figure 12: Non-dimensional bubble radius plot for case 1 of Liquid Nitrogen
Figure 13: Actual bubble radius plot for case 1 of Liquid Nitrogen
From the figure 12,13, it is evident that the non-dimensional bubble radius grows along the converging section from $1$ to $3.4$ up to the throat($x=0$ indicates the beginning of the converging section as per Brennen Area profile,$x=1$ indicates the throat section (or) beginning of the diverging section and $x=2$ indicates the exit of diverging section), and dampens from the throat to the initial value of 1 at the exit of the nozzle and the region away from the nozzle exit.
Figure 14: Non dimensional Void fraction plot for the case1 of Liquid nitrogen
The void fraction plot, as shown in figure 14, increases continuously up to 3.8 ×$\times 10^{-4}$ after the nozzle exits from the location of 2.7 to the far downstream of the nozzle to reach the value of 1 which corresponds to fully vapor condition. From the plot of Coefficient of Pressure given in the figure 15, the same behaviour is inferred.
Figure 15: Coefficient of Pressure plot for the case 1 of Liquid nitrogen
The magnitude of Pressure and Thermal source term is shown as follows, which shows that the magnitude of the thermal term for liquid nitrogen is 10 times the same for water,
Figure 16: Magnified view of coefficient of Pressure plot for case 1 of Liquid nitrogen
The magnitude of Pressure and Thermal source term is shown as follows, which shows that the magnitude of the thermal term for liquid nitrogen is 10 times the same for water, indicating the domination of thermal term in cryogenic liquids compared with isothermal fluids such as water 17. The following figures 17, 18 shows the source terms plotted for case1
Figure 17: Magnitude of Pressure and Thermal source term plot
Figure 18: Magnified view of pressure source term
2.5.2 Case-2
Plot for the same are predicted for an Initial void fraction of $\alpha_{0}=3.1\times 10^{-6}$ and coefficient of pressure $C_{p}$=0
Figure 19: Non-dimensional bubble radius plot for the case2 of Liquid Nitrogen
Figure 20: Actual bubble radius plot for case 2 of Liquid Nitrogen
Figure 21: Non-dimensional Void fraction plot for the case 2 of Liquid nitrogen
Figure 22: Coefficient of Pressure plot for the case 2 of Liquid nitrogen
Figure 23: Magnitude of Pressure and Thermal source term plot
Figure 24: Magnified view of pressure source term
2.5.3 Case-3
Plot for the same is predicted for an Initial void fraction of $\alpha_{0}=3.0\times 10^{-6}$ and coefficient of pressure $C_{p}$ = 0
Figure 25: Non-dimensional bubble radius plot for the case 3 of Liquid Nitrogen
Figure 26: Actual bubble radius plot for case 3 of Liquid Nitrogen
Figure 27: Non-dimensional Void fraction plot for the case 3 of Liquid nitrogen
Figure 28: Coefficient of Pressure plot for the case 3 of Liquid nitrogen
Figure 29: Magnitude of Pressure and Thermal source term plot
It is concluded that the 1D numerical model for cavitating flows with the thermal effects resulted in successful benchmarking of the Wang &\&& Brennen [9] solution for water. The extension of the same model for cryogenic cavitating nozzle flows resulted in the damping nature of solutions as the behavior of cryogenic liquid nitrogen distinctly differs from the water. Overall, it can be ascertained that the 1D numerical model gives a limited predictability of the nature of cavitating flows, which are restricted to Bubble Radius, Void fractions, Coefficient of Pressure, and the thermal and pressure source terms corresponding to a single bubble as the governing Rayleigh Plesset equation itself is valid for dynamics of the single bubble only. However, cavitation is a cluster phenomenon that demands the extension of the studies to two-dimensional numerical and experimental studies, whose attempted results are given in the subsequent sections.
3 Two-dimensional model CFD Studies
The following section systematically starts with an overview of the methodology adopted for 2D numerical studies using the commercial CFD solver ANSYS Fluent. The simulations have been carried out to capture the cavitating flows in a Venturi profile chosen to be symmetric, with a turbulence model to get insight into the actual experimental work. Simulations were done for cryogenic liquid nitrogen. The Governing Equations, such as continuity and momentum for the multiphase model, are solved using finite volume fluent code. The current chapter discusses the viscous modeling, interfacial effects, solution methodology, and the boundary conditions used for the benchmarking test case of Hord et al. [5] and the extension of the same for current 2D numerical work.
3.1 Turbulence Modelling
The realistic cases of cavitating nozzle and venturi flows are turbulent in nature. Hence, a standard turbulence model has been incorporated into the simulations. Realizable $k-\varepsilon$ was used in simulations. The chosen model of $k-\varepsilon$ is in line with the internal flows, and the coefficients for the same are the standard values built into the ANSYS-fluent. The model has two transport equations, one for turbulent kinetic energy k and another for turbulent dissipation $\varepsilon$. The two transport equations are given as follows
$\frac{\partial}{\partial t}(\rho k)+\frac{\partial}{\partial x_{i}}(\rho ku_{i})=\frac{\partial}{\partial x_{j}}\bigg{[}\bigg{(}\mu+\frac{\mu_{t}}{\sigma_{k}}\bigg{)}\frac{\partial k}{\partial x_{j}}\bigg{]}+G_{k}+G_{b}-\rho\varepsilon-Y_{m}$ (17)
And
$\frac{\partial}{\partial t}(\rho\varepsilon)+\frac{\partial}{\partial x_{i}}(\rho\varepsilon u_{i})=\frac{\partial}{\partial x_{j}}\bigg{[}\bigg{(}\mu+\frac{\mu_{t}}{\sigma_{\varepsilon}}\bigg{)}\frac{\partial\varepsilon}{\partial x_{j}}\bigg{]}+\rho C_{1}S_{l}\varepsilon+C_{1\varepsilon}\frac{\varepsilon}{k}(G_{k}+C_{3\varepsilon b}G_{b})\\ -C_{2\varepsilon}\rho\frac{\varepsilon^{2}}{k+\sqrt{\nu\varepsilon}}$ (18)
Where, C1=max $\bigg{[}0.43+\frac{\eta}{\eta+5}\bigg{]}$ , $\eta=\frac{S_{l}k}{\varepsilon}$, $S_{l}=\sqrt{2S_{ij}S_{ij}}$ is the modulus of mean strain rate tensor.
The turbulent viscosity $\mu_{t}=\rho C_{\mu}\frac{k^{2}}{\varepsilon}$, where $C_{\mu}$ is a function of mean strain rate and mean rotation rate, the angular velocity of the system rotation, and the turbulence fields such as k and $\varepsilon$. The model constants are given by,
$C_{1\varepsilon}=1.44,\leavevmode\nobreak\ \leavevmode\nobreak\ C_{2\varepsilon}=1.9,\leavevmode\nobreak\ \leavevmode\nobreak\ \sigma_{k}=1,\leavevmode\nobreak\ \leavevmode\nobreak\ \sigma_{\varepsilon}=1.2$
The term Gk indicates the production of turbulent kinetic energy and is given by,
$G_{k}=\mu_{t}S_{l}^{2}$
The turbulence production due to buoyancy forces can be computed from the following relation
$G_{b}=\beta_{g}\frac{\mu_{t}}{Pr_{t}}\frac{\partial T}{\partial x_{i}}$
3.2 Modelling Interfacial effects
3.2.1 Schnerr Sauer model for Mass Transfer
Estimation of mass transfer between the two phases has to be taken care of for complete cavitation modelling. The Schnerr and Sauer model addresses the net mass transfer from the liquid to the vapour. The vapour transport equation is given by
$\frac{\partial}{\partial t}(\alpha\rho_{\nu})+\nabla.(\alpha\rho_{\nu}\bar{\nu})=\frac{\rho_{\nu}\rho_{l}}{\rho}\frac{D\alpha}{Dt}$
The net source term for mass is given by the right side of the above equation as follows
$\dot{m}=\frac{\rho_{\nu}\rho_{l}}{\rho}\frac{D\alpha}{Dt}$
Schnerr-Sauer model incorporated an expression that relates the vapour volume fraction to the bubble number per unit volume of liquid, which is defined as
$\alpha=\frac{\eta_{b}\frac{4}{3}\pi R_{b}^{3}}{1+\eta_{b}\frac{4}{3}\pi R_{b}^{3}}$
Rb is the bubble radius, and $\eta_{b}$ is the bubble population number per unit volume of the liquid. In this model assumes that the number of bubbles is neither created nor destroyed. To model bubble dynamics, the Rayleigh-Plesset equation in simple form was incorporated that was built with Schnerr-Sauer model, which is given by
$\frac{D\alpha}{Dt}=\sqrt{\frac{(P_{b}-P)^{2}}{\rho_{l}^{3}}}$ (19)
Where Pb is the bubble pressure, P is the pressure which is felt far away from the bubble. Using the above relations, the mass source term is
$\dot{m}=\frac{\rho_{\nu}\rho_{l}}{\rho}\alpha(1-\alpha)\frac{3}{R_{b}}\sqrt{\frac{(P_{b}-P)^{2}}{\rho_{l}^{3}}}$
3.3 Interfacial Heat Transfer Effects
Modelling the convective heat transfer coefficient $h_{b}$ was given by Ranz and Marshall [14] model, which is inbuilt in Fluent and is given as follows
$h_{b}=h_{pq}=\frac{6k_{q}\alpha_{p}\alpha_{q}Nu_{p}}{d_{p}^{2}}$
Here kq is the thermal conductivity of the qth face. The Nusselt number was determined from
$Nu_{p}=2.0+0.6Re_{p}^{1/2}Pr^{1/3}$
Where Rep is the Reynolds number of the pth phase and Pr is the Prandtl number of the qth phase.
3.4 Benchmark of 2D simulations with experimental work of Hord.et al
An attempt was made to benchmark the 2D numerical work involved in this work with the experimental test case of Hord 121B [5] for validating the cavitating length for liquid hydrogen flows in a venturi profile.
A finite volume mesh was done using Ansys for the actual profile used by Hord for his experiment. A grid independence study was done for the Venturi Profile, and the grid independent solution was obtained for a mesh with 51,335 nodes and 50527 elements. The following figure shows the meshed model of the venturi profile.
Figure 30: Mesh details of Hord experimental venturi profile
The simulation was done with the boundary conditions, which were taken from the experimental data of Hord et al. [5].
3.4.1 Solution Methodology
The simulations were performed using fluent with the Eulerian Multiphase model. An implicit finite volume scheme is based on a segregated pressure-based solver with SIMPLE as the pressure-velocity coupling. A second-order upwind scheme was used for continuity, momentum, energy, turbulent dissipation, and Kinetic energy.
Being turbulent flow, the wall y$+$ was assumed to be 30, and the resolution of grid element size near the boundary was 2 $\times$ 10-04 m. A default standard wall function was chosen for solution methodology [8].
The residuals were set to be 0.001 for continuity, velocity, K, and $\varepsilon$. For Energy and volume fraction, the residuals were set to be 1-06.
Following are the Simulation results for the Contour of Vapour fraction obtained using Ansys fluent
3.4.2 Simulation Results
The following table31 is the actual experimental result of Hord for various test cases from the flow visualization study.
Figure 31: Table showing experimental test cases of Hord
The following figure32 shows the simulation results of Hord test case 121B. The blue fill indicates the complete liquid phase, and the red fill indicates the complete vapor phase.
Figure 32: Contour of volume fraction from the simulation results of Hord profile
Figure 33: Plot for volume fraction as a function profile length
From the vapour volume fraction plot33, it is evident that the vapour phase dominates in the cavitating regime, and the plot shows the cavitating length of 0.085 m or 8.5 cm.
The simulation results for other test cases are as follows:
Figure 34: Contour of volume fraction for test case 124B
Figure 35: Contour of volume fraction for test case 126A
Figure 36: Contour of volume fraction for test case 128A
3.4.3 Estimating Cavitating length
The details of the measured Cavitating length from the contour of the void fraction is shown in the below figure37
Figure 37: Cavitation length measurement details
[…]
Error prediction in cavitating length against the actual experimental length for different test cases are shown in the figure38 containing the table
Figure 38: Table showing error in the simulation results of Hord
Similar work was done by Rodio et al. [8], and the contour of the void fraction of the vapor obtained for the Hord Test case 121B [5] is shown in the following figure.
Figure 39: Contours of the void fraction obtained by Rodio et.al [8] for Hord test case 121B [5]
It was observed from the results that the simulation performed by Rodio et al. [8] overpredicted the experimental cavitating length observed by Hord et al. [5]. The following measurement proofs validate the above interpretation of the result.
Figure 40: Cavitation Length measurement details for Rodio et.al [8] work
The calculated cavitating length from the above contours of volume fraction gives a length of 11.8235 cm against the actual Experimental cavitating length of 8.26 cm, measured by Hord et al. [5]. However, the wrong prediction of cavitation length may not be precisely due to the model used, as mesh methods and wall functions have a great influence on solution convergence. Having benchmarked the 2D numerical model for cryogenic cavitating venturi with Hord et al. [5] experimental test cases for predicting the cavitating length in the venturi profile. The same is now extended for the current work, the details of which are given in the following section.
4 2D simulation for Current Work
4.1 Computational Model Details
A steady state, two-dimensional, incompressible, turbulent Eulerian mixture model was developed based on finite volume formulation. The details of the computational model are shown below
Figure 41: Computation model with domain or boundary
The figure41depicts the actual 2D model of the venturi region of the test section developed for simulations. The dimensions correspond to the actual dimensions of the experimental setup. The boundaries include an inlet, wall, symmetry, and an outlet.
Domain discretization or meshing was done with Ansys and Pointwise. The entire domain was discretized using structural quadrilateral elements. The wall y+ requirement for $k-\varepsilon$ model was assumed to be around 30 and accordingly the thickness of the first computational grid was fixed.
4.2 Boundary conditions
The entire domain was divided into three regions: inlet, outlet, and walls. Only the inlet and outlet boundary conditions were found to be very important as they play a very significant role in the convergence of solutions.
The inlet and outlet pressure conditions from the experiments were used as boundary conditions, with the inlet temperature assumed as 77 K for the liquid phase of liquid nitrogen. Also, the inlet vapour fraction is given 0 as the inlet condition is fully liquid.
Adiabatic wall boundary conditions were given as there were no assumptions of heat transfer effects between the wall boundary and fluid.
The outlet boundary condition was atmospheric pressure, as the experiment’s venturi profile opens to the atmosphere. Liquid temperature condition at the outlet is assumed to be 77 K, and vapour fraction at the outlet is assumed to be 0 as the single-phase liquid was desired.
4.3 Solution Methodology
The same solution methodology was used for the Hord benchmark simulation. It is the exact segregated pressure-based solver with SIMPLE as the pressure-velocity coupling. A Second Order Upwind scheme was used for continuity, momentum, energy, turbulent dissipation, and kinetic energy.
Under-relaxation parameters for vapourization mass were set to 0.5, and the volume fraction was set to 0.95.
The residuals were set to be 0.001 for continuity, velocity, K, and $\varepsilon$. For Energy and volume fraction, the residuals were set to be 1 e-06. The simulation results for the contour of vapour fraction obtained using Ansys fluent are discussed in following section.
4.4 Grid Independence
A grid independence study was carried out to find the optimum number of grid points that ensures correct capturing of flow characteristics and ensures that the solution does not change with further increase in the number of grid points.
The domain was discretized using finite volume meshing in Ansys with grid elements of 8346, 40926, and 102482 elements. It was found that a grid-independent solution was obtained with a mesh of 40926 grid elements. The following picture shows the meshing details. For near-wall treatment, a default wall y+ $>$ 30 for turbulent flow was assumed, and accordingly, the size of the first grid element near the wall was chosen for meshing. The following picture shows the meshing details
Figure 42: Mesh detail for the venturi
4.5 2D Numerical results
This section discusses the simulation results of the symmetric venturi corresponding to 40926 grid elements. The Boundary conditions used in the simulations were in line with experimental observations.
4.5.1 Boundary Conditions
The following table 43 gives the details of the experimental observations. The same values of pressure are used for 2D simulations as boundary conditions.
Figure 43: Table showing experimental data
The simulation results obtained from Ansys-Fluent for various test cases are as follows.
4.5.2 Contour of Volume fraction(Liquid)
Figure 44: Contours of volume fraction for test case 396
Figure 45: Contours of volume fraction for test case 397
Figure 46: Contours of volume fraction for test case 398
Figure 47: Contours of volume fraction for test case 399
The blue fill indicates a complete vapour phase, and the red fill indicates a fully liquid phase. The yellow region corresponds to a vapour fraction of 0.6, which indicates a thin two-phase regime, which is a point where the vapor is ready to convert back to the single-phase liquid. It is evident from the above figures that there is a distinct cavitation regime from the throat section up to a certain length in the downstream section of the venturi that corresponds to the Cavitating length of the Venturi.
4.5.3 Cavitating Length Measurement
The sample calculation of the cavitating length is shown as follows
Figure 48: Sample cavitating length measurement for symmetric venturi
From the above calculation, the cavitating length was found by subtracting the x-coordinate at the end of the vapour regime from the x-coordinate at the throat location, which results in the cavitating length of 45.668 mm or 4.56 cm. A similar calculation is done for other cases, and the error in the cavitating length of simulation against the experiments is shown in the following table 49
Figure 49: Table showing the error in simulation cavitating length
[…]
5 Experimental Observations
This section puts forth the experimental observations of two experiments, one with the acrylic venturi test section and the other with the aluminum venturi test section. The acrylic venturi was incompatible with the cryogenic temperature flow conditions, resulting in the crack of the test section itself.
Figure 50: Schematic of the Experimental setup.
Figure 51: Experimental setup.
Following this, an attempt was made using the same experiment with the aluminium venturi. The flow visualization images during the cavitating time of venturi are presented, and an attempt was made to evaluate the experimental cavitation length from these visualization images.
5.1 Observations from experiment - 1
The experiment was carried out with an initial Dewar pressure of 2.5 bar, and the outlet of the venturi test section was opened to atmospheric pressure. After an initial transient time for chilling down of the set-up, the developed flow in the test section appeared 4.30 mins from the start of the experiment. Immediately, the nature of the flow of liquid nitrogen turned stratified while flowing inside the test section. This was because of the crack, which developed in the adapter region upstream of the test section, which had a metal-to-acrylic contact. Also, it was observed that the acrylic test section, in due time, resulted in multiple cracks at different spots, which can be seen from the following figure, which was taken from the 42547th frame record of the high-speed camera.
Figure 52: Test section of the experiment - 1
In the above experiment, there was no sign of any developed single-phase flow of liquid nitrogen, and there was no recognized cavitating regime because of the stratified nature of the flow, which was believed to be caused by the cracks in the test section. So, the image data obtained did not give any physical insight into the cavitating nature of the flow and the length. However, the pressure measurement readings closely matched the cavitating regime’s operating conditions.
5.2 Observations from experiment - 2
The experiment was carried out with an initial Dewar pressure of 1.2 bar, and the outlet of the aluminium venturi test section was opened to the atmosphere. The time taken for chilling down was reduced compared to the previous experiment, as the entire test section was wound with foam insulation. The developed flow in the test section appeared at 2 min from the start of the experiment. The image was captured at 198th sample scan of the data acquisition system from the video taken by the high-speed camera.
Figure 53: Flow visualization image captured during experiment - 2
The figure 53 shows a distinct vapor regime, which starts from the throat region and extends to some regions in the diverging section of the venturi. The reading from the pressure measurement indicated a near-cavitation operating condition.
Figure 54: Sample cavitating length measurement
The figure 53 was image processed to get the cavitation length, shown in the figure 54.
The following figure 55 shows the images and corresponding cavitating length for the other test cases.
Figure 55: Figure showing various test cases
6 Conclusion
  1. 1.
    The attempted numerical 1D model based on the extension of Wang and Brennen [9] model with an additional thermal term in line with numerical 1D model of Rodio [8] was implemented using MATLAB code. Following are the conclusions drawn from the current 1D numerical studies:
    • A benchmark study of this model with water was done, and the results matched well with Wang and Brennen [9] model. The superimposed plots of the current model with the actual plot of Wang and Brennen [9] are shown in the figures4,5,6.
    • It was also interpreted that the thermal term in a cavitating water flows had a negligible role to play in bubble growth or collapse as the plots for the magnitude of source terms 10 &11had an order of 10-1 for the thermal term and 104 for pressure term.
    • However, the present model has shortcomings in predicting realistic performance parameters of cavitating venturi.
  2. 2.
    The 2D numerical model was implemented in ANSYSFluent, and the following are the outcomes of the simulation studies.
    • The cavitating length predicted for benchmarking the experimental results of hydrogen cavitation for different test cases of Hord et al. [5] whose details are shown in the figures 37,38 for validating the developed 2D numerical model.
    • A similar procedure was extended for the actual symmetric venturi using liquid nitrogen using the experimental data as the boundary conditions as shown in the figure 43, and the cavitating length and errors for the same with experimental cavitating lengths were predicted as detailed in the figures48 &49, which showed an error of less than $5\%$ that could be accepted as reasonable predictions.
    The following are the outcomes from the Experimental observations
    • The acrylic venturi test section did not withstand the low-temperature liquid nitrogen flow, resulting in leaks near the upstream venturi and altering the flow pattern to stratified conditions as shown in 52. Hence, a distinct cavitating flow could not be observed in experiment 1.
    • The experiment-2, which employed an aluminum venturi, showed distinct cavitating regimes when operated at an initial dewar pressure of 1.5 bar as shown by the figure 53
    • The cavitating length measured from the experimental image is shown in the figures54 & 55
Acknowledgement We acknowledge Prof.Kannan Iyer and Dr.Nandhakumar (LPSC) for their active discussions. Mr.Dinesh and Mr.Bipin of Thermal Engineering lab are recognized for their assistance with Experiments.
Appendix A: Numerical Method for Modelling 1D Cavitating nozzle / Venturi flow using Matlab
Figure 56: steps involved in numerical coding
Step1: List of Governing Equations and closures:
Continuity and Momentum equation
$\dfrac{\partial(1-\alpha)A}{\partial t}+\dfrac{\partial(1-\alpha)Au}{\partial x}=0,$
$\dfrac{\partial u}{\partial t}+u\dfrac{\partial u}{\partial x}=-\dfrac{1}{2(1-\alpha}\dfrac{\partial C_{p}}{\partial x}$
Pressure Coefficient
$C_{p}(x,t)=\dfrac{P(x,t)-P_{0}}{0.5\rho_{l}u_{0}^{2}}$
Void fraction equation
$\alpha(x,t)=\dfrac{\dfrac{4}{3}\pi\eta R(x,t)^{3}}{\left(1+\dfrac{4}{3}\pi\eta R(x,t)^{3}\right)}$
Modified Rayleigh-Plesset equation
$\left[R\ddot{R}+\dfrac{3}{2}\dot{R}^{2}\right]+\dfrac{2S}{\rho_{L}R}+\dfrac{4\mu}{\rho_{L}}\dfrac{\dot{R}}{R}=\dfrac{P_{g0}}{\rho_{L}}\left[\dfrac{R_{0}}{R}\right]^{3\gamma}+\dfrac{\left[P_{\nu}(T_{\infty})-P_{\infty}(t)\right]}{\rho_{L}}-\dfrac{dP_{\nu}}{dT}\dfrac{\rho_{\nu}L}{\rho_{L}}\dfrac{\dot{R}}{h}$
Ranz &\&& Marshall convective heat transfer model
$h_{b}=\dfrac{Nu_{b}K_{l}}{2R}$
$Nu_{b}=2+0.6Re_{b}^{\frac{1}{2}}Pr^{\frac{1}{3}}$
$Re_{b}=2R|\nu-u|/\lambda_{l}$
$Pr=\dfrac{C_{pl}\mu_{l}}{K_{l}}$
Bubble momentum equation
$\rho_{\nu}\dfrac{D\nu}{Dt}+\dfrac{1}{2}\rho_{l}\left(\dfrac{D\nu}{Dt}-\dfrac{Du}{Dt}\right)=-\dfrac{\partial P(x,t)}{\partial x}-\dfrac{3}{8}\rho_{l}C_{D}\times\dfrac{(\nu-u)|\nu-u|}{R}$
Step2: Non-dimensionalizing the Equations:
All the equations mentioned above were non-dimensionalized by using following dimensionless variables
$\overline{u}=\dfrac{u}{u_{0}};\overline{x}=\dfrac{x}{R_{0}};\overline{R}=\dfrac{R}{R_{0}};\overline{t}=\dfrac{tu_{0}}{R_{0}};\overline{v}=\dfrac{v}{u_{0}};\overline{\eta}=\dfrac{\eta}{R_{0}^{3}};\overline{L}=\dfrac{L}{R_{0}}$
where the subscript ”0” corresponds to the upstream value. Some of the non-dimensional numbers, such as
$Re=\dfrac{\rho_{L}uR}{\mu_{E}};We=\dfrac{\rho_{L}uR}{S};\sigma=\dfrac{P-P_{v}}{\dfrac{1}{2}\rho_{L}U^{2}}$
After applying these dimensionless terms to the actual steady case equation, the final system of dimensionless equations is of the form
$\left(1-\alpha\right)\overline{u}\overline{A}=\left(1-\alpha_{0}\right)=const$
$\overline{u}\dfrac{d\overline{u}}{d\overline{x}}=-\dfrac{1}{2\left(1-\alpha\right)}\dfrac{dc_{p}}{d\overline{x}}$
$\overline{R}\left(\overline{u}^{2}\dfrac{d^{2}R}{dx^{2}}+\overline{u}\dfrac{d\overline{u}}{d\overline{x}}\dfrac{d\overline{R}}{d\overline{x}}\right)+\dfrac{3}{2}\overline{u}^{2}\left(\dfrac{d\overline{R}}{d\overline{x}}\right)^{2}+\dfrac{4\overline{u}}{Re\overline{R}}\dfrac{d\overline{R}}{d\overline{x}}+\dfrac{2}{We}\left(\dfrac{1}{\overline{R}}-\dfrac{1}{\overline{R}^{3\gamma}}\right)=-\dfrac{C_{p}}{2}-\dfrac{\sigma}{2}\left(1-\dfrac{1}{\overline{R}^{3\gamma}}\right)\\ -\dfrac{dP_{v}}{d\overline{T}}\dfrac{L_{ev}\rho_{v}}{\rho_{l}h_{b}}\overline{u}\dfrac{d\overline{R}}{d\overline{x}}$ (20)
$\alpha(x,t)=\dfrac{\dfrac{4}{3}\pi\overline{\eta}\overline{R}^{3}}{\left(1+\dfrac{4}{3}\pi\overline{\eta}\overline{R}^{3}\right)}$
$\overline{v}\dfrac{\partial\overline{v}}{\partial\overline{x}}=\left\{-\dfrac{1}{\left(\rho_{v}+\dfrac{1}{2}\rho_{l}\right)\overline{v}u_{0}^{2}}\right\}\left\{\dfrac{\partial P}{\partial\overline{x}}-\dfrac{1}{2}\rho_{l}u_{0}^{2}\overline{u}+\dfrac{3}{8}\dfrac{C_{D}\rho_{l}}{\overline{R}}u_{0}^{2}|\left.\left(\overline{v}-\overline{u}\right)\right|\right\}$ (21)
The above-governing equations were integrated numerically using a higher-order adaptive Runge-Kutta scheme known as the Dormand-Prince scheme. Details of the adaptive numerical methods are discussed in detail in the next section.
Step3: Numerical solution procedure:
  • Dormand-Prince is called adaptive as it attempts to optimize the interval size and reduce the computational cost. The adaptive techniques algorithm should know when the interval (h) size must be adjusted.
  • Adaptive methods work by comparing the solution of an Ordinary Differential Equation by two methods, say Euler &\&& Heun’s methods[22]. From the solution difference, we find the relative error ofthe Euler method with respect to Heun’s method (as Heun’s method is better in comparison with Euler)[22].
  • Then, this Error is compared with the allowable error in our computation, and based on that, a scaling factor is found for updating the interval size at each step.
General procedure for finding Scaling parameter for ‘h’ with an illustrated example
For a given initial value problem,
$y^{(1)}(t)=f(t,y(t))$
$y(t_{0})=y_{0}$
Step1: Finding the solution using any two methods of numerical integration. Here, in this case, let the solution be done using Euler and Heun’s method
$K_{1}=f(t_{k},y_{k})$
$K_{2}=f(t_{k}+h,y_{k}+hK_{1})$
$y_{tmp}=y_{k}+hK_{1}$
$z_{tmp}=y_{k}+h\dfrac{K_{1}+K_{2}}{2}$
Where $y_{tmp}$ and $z_{tmp}$ are the solutions obtained from Euler and Heun’s methods, respectively.
Step2: The difference between the two solutions, $|y_{tmp}-z_{tmp}|$ is the Error. This error obtained is the error in the Euler method compared to Heun’s method, and the error is of order $O(h^{2})$. Therefore, error can be written as a function of $h^{2}$
$|y_{tmp}-z_{tmp}|=Ch^{2}$ (22)
Step3: Substituting scaling factor. Scaling ’h’ by some factor ’s’
$|y_{tmp}-z_{tmp}|=Ch^{2}=C(sh^{2})$ (23)
Step4: Comparing the Error with allowable error &\&& finding ’s’
$C(sh^{2})<\varepsilon_{abs}$
The contribution of the maximum error at the kth step should be proportional to the width of the interval relative to the whole interval
$C(sh)^{2}<\varepsilon_{abs}\dfrac{sh}{t_{f}-t_{0}}$
$C(sh)^{2}=\dfrac{1}{2}\varepsilon_{abs}\dfrac{sh}{t_{f}-t_{0}}=\dfrac{\varepsilon_{abs}sh}{2\left(t_{f}-t_{0}\right)}$ (24)
$\Rightarrow s(Ch^{2})=\dfrac{\varepsilon_{abs}h}{2\left(t_{f}-t_{0}\right)}$ (25)
Putting equation 22 in 24
$s|y_{tmp}-z_{tmp}|=\dfrac{\varepsilon_{abs}h}{2\left(t_{f}-t_{0}\right)}$
$\Rightarrow s=\dfrac{\varepsilon_{abs}h}{2\left(t_{f}-t_{0}\right)|y_{tmp}-z_{tmp}|}$ (26)
Updating ’h’,
$h^{\ast}=s\times h$
where, $h^{\ast}$ is the updated value of h.
Depending on the value of ’s’, the value of ’h’ is updated as follows
  • If s $\geq$ 2, $\qquad h^{\ast}=2\times h$
  • If 1 $\leq$ s $<$ 2, $\qquad h^{\ast}=h$
  • If s $<$ 1, $\qquad h^{\ast}=\dfrac{h}{2}$ (and try again)
Algorithm for Dormand Prince Adaptive method
In the Dormand Prince adaptive method, RK4 and RK5 methods are compared generally
Step1: Finding the slopes $k_{1}$ to $k_{7}$.
$k_{1}=hf\left(t_{k},y_{k}\right)$
$k_{2}=hf\left(t_{k}+\dfrac{1}{5}h,y_{k}+\dfrac{1}{5}k_{1}\right)$
$k_{3}=hf\left(t_{k}+\dfrac{3}{10}h,y_{k}+\dfrac{3}{40}k_{1}+\dfrac{9}{40}k_{2}\right)$
$k_{4}=hf\left(t_{k}+\dfrac{4}{5}h,y_{k}+\dfrac{44}{45}k_{1}-\dfrac{56}{15}k_{2}+\dfrac{32}{9}k_{3}\right)$
$k_{5}=hf\left(t_{k}+\dfrac{8}{9}h,y_{k}+\dfrac{19372}{6561}k_{1}-\dfrac{25360}{2187}k_{2}+\dfrac{64448}{6561}k_{3}-\dfrac{212}{729}k_{4}\right)$
$k_{6}=hf\left(t_{k}+h,y_{k}+\dfrac{9017}{3168}k_{1}-\dfrac{355}{33}k_{2}-\dfrac{46732}{5247}k_{3}+\dfrac{49}{176}k_{4}-\dfrac{5103}{18656}k_{5}\right)$
$k_{7}=hf\left(t_{k}+h,y_{k}+\dfrac{35}{384}k_{1}+\dfrac{500}{1113}k_{3}+\dfrac{125}{192}k_{4}-\dfrac{2187}{6784}k_{5}+\dfrac{11}{84}k_{6}\right)$
Step2: Finding the solution using RK4 and RK5 method.
The solution by Runge-Kutta 4th order method is given by:
$y_{k+1}=y_{k}+\dfrac{35}{384}k_{1}+\dfrac{500}{1113}k_{3}+\dfrac{125}{192}k_{4}-\dfrac{2187}{6784}k_{5}+\dfrac{11}{84}k_{6}$
The solution by Runge-Kutta 5th order method is given by:
$y_{k+1}=y_{k}+\dfrac{5179}{57600}k_{1}+\dfrac{7571}{16695}k_{3}+\dfrac{393}{640}k_{4}-\dfrac{92097}{339200}k_{5}+\dfrac{187}{2100}k_{6}+\dfrac{1}{40}k_{7}$
Step3: The Error in the RK4 method is calculated by finding the difference between the two solutions. The error for the Dormand Prince method is as shown below:
$\left|z_{k+1}-y_{k+1}\right|=\left|\dfrac{71}{57600}k_{1}-\dfrac{71}{16695}k_{3}+\dfrac{71}{1920}k_{4}-\dfrac{17253}{339200}k_{5}+\dfrac{22}{525}k_{6}-\dfrac{1}{40}k_{7}\right|$
Step4: Finding the scaling factor and optimum interval size. The scaling factor for Dormand Prince is given by
$s=\left(\dfrac{\varepsilon h}{2\left|z_{k+1}-y_{k+1}\right|}\right)^{\dfrac{1}{5}}$
The optimum interval size is given by Error prediction methods by Shampine[22]
$h_{opt}=sh$
The above algorithm is incorporated in Matlab coding, and the model was solved for cases involving water and liquid nitrogen. Numerical solutions with water after including the thermal model are discussed in the following section.